OpenStudy (anonymous):

FIRST CORRECT ANSWER GETS A MEDAL! Kyle has a total of 16 coins in his pocket. All the coins are nickels or quarters, and they have a total value of $2.20. How many nickels does Kyle have?

OpenStudy (gorv):

let no of nickel coin = x than quarter = 16-x

OpenStudy (gorv):

value of nickel * x+value of quarter *(16-x)=2.20

OpenStudy (gorv):

put the values of coins here

OpenStudy (anonymous):

I still don't understand

OpenStudy (anonymous):

9 nickles

OpenStudy (anonymous):

nickels*

OpenStudy (anonymous):

it's not in the multiple choices

OpenStudy (gorv):

ok what is the value of nickel??

OpenStudy (anonymous):

$2.20?

OpenStudy (gorv):

nickel and quarter are there

OpenStudy (gorv):

nickel worth???

OpenStudy (anonymous):

lol. If you have 2.20 and 9 of the 16 coins are nickels. You would do 9x5

OpenStudy (anonymous):

wait never mind i got it. thank you

OpenStudy (gorv):

nickel equal to 5 cent

OpenStudy (gorv):

quarter = 25 cent

OpenStudy (gorv):

we total have 16 coin now let we x nickel so remaining coins are quarter that is =16-x

OpenStudy (gorv):

value of nickel coin= 5cent *no of nickel coin=5*x value of quarter=25*(16-x)

OpenStudy (gorv):

2.2 dollar=2.2*100 =220 cent and sum of their value = 2.2 dollar=220 cent so 5x+25*(16-x)=220

OpenStudy (gorv):

solve it now